# Glitch in Time: Exploiting Temporal Misalignment of IMU For Eavesdropping

### v2

@@ Abstract.
copyright: none
@@ 2.1.4. Spoken Language Understanding (SLU):
Paper Sensor Recognition
Evaluation
Metric
Accuracy
@ 200 Hz
@@
PitchIn (Han et al., 2017) Acc
Words
Accuracy NA
Spearphone (Anand et al., 2021) Acc
Digits + Words
F-measure NA
AccelEve (Ba et al., 2020) Acc
Digits + Letters
Accuracy NA
@@
InertiEar (Gao et al., 2022) Acc+Gyro
Digits + Letters
Accuracy 78.8%
StealthyIMU (Sun et al., 2023) Acc
Spoken Language Understanding
@@
STAG [Ours] Acc+Gyro
Spoken Language Understanding
@@ 2.2. Related Works
AccelEve (Ba et al., 2020) marks the first study exploiting loudspeaker proximity to Inertial Measurement Units (IMUs). It employs a deep learning model (DenseNet) to reconstruct speech from accelerometer signal spectrograms, achieving a top-1 accuracy of 78% for digits and 55% for digits plus letters. Spearphone (Anand et al., 2021) utilizes altered accelerometer readings and machine learning techniques to classify gender with over 90% accuracy and identify speakers with over 80% accuracy. Additionally, Spearphone can reconstruct speech with a word detection accuracy of 67%.
@@ 4.1. Temporal misalignment
However, this method encountered a significant limitation. Despite their staggered start times, both accelerometer instances produced readings with identical timestamps. This outcome was due to both instances drawing data from the same physical accelerometer sensor. As a result, the sensor values, which are dependent on the changes in the sensor’s state, were the same for both instances. Moreover, the Android operating system does not guarantee a fixed 200 Hz sampling rate; the actual rate can fluctuate above or below this figure. This rate flexibility allows Android to provide the most recent sensor values to any software instance of the accelerometer. Due to this system behavior and the shared use of the physical sensor, both software instances ended up capturing essentially identical data, as reflected in their synchronized timestamps.
Lesson Learned: Attempting to create multiple software instances of the accelerometer within an Android application to simulate a higher sampling rate is ineffective. The Android operating system’s handling of sensor data results in identical timestamps for readings from both instances, due to their shared use of the same physical sensor. This highlights the limitations imposed by the OS and hardware design on sensor data manipulation.
@@
The cause of this synchronized timestamp phenomenon can be traced back to two main factors. Firstly, Android’s management of sensor sampling rates is inherently flexible; while a maximum rate is specifiable, the actual sampling frequency can vary and is adjusted in real-time by the operating system. This variability in the sampling intervals could result in the gyroscope and accelerometer inadvertently aligning their readings, despite our efforts to stagger their activation. Secondly, the design of IMU sensors, which encompasses both accelerometers and gyroscopes, is inherently inclined towards synchronized functioning. These sensors are typically integrated at both hardware and software levels to facilitate efficient and precise motion detection, which naturally leads to their simultaneous data sampling and reporting. A key finding was the consistent zero variance in sampling delays, indicating potential for precise gyroscope sample alignment in future experiments.
Lesson Learned: Introducing gyroscope and accelerometer sensors with a staggered initialization also fails to achieve desynchronized readings, as the Android OS manages sensor sampling rates flexibly and IMU sensors are designed to operate synchronously. This experiment underlines the challenge of achieving temporal misalignment through software alone, due to the inherent synchronized functioning of IMU components and the OS’s real-time adjustments to sampling rates.
@@
4.1.3. Scenario 3: introducing magnetometer.
The distinct hardware configurations of sensors, particularly the STMicroelectronics LSM6DSM IMU (ST, 2024) found in the Samsung A31/A33 and the Bosch BMI160 IMU (BOSCH, 2024) used in the Pixel 5, involve direct connections between the IMU and the host, with the magnetometer interfacing through the IMU. This architectural design involves a FIFO buffer within the IMU that aids in efficient high-frequency data capture and energy savings. However, it can introduce synchronization problems when integrating a Magnetometer which does not have its own FIFO buffer and whose values are stored in the FIFO of IMU—leading to data misalignments, as our research indicates.
Lesson Learned: Integrating a magnetometer with gyroscope and accelerometer systems can effectively introduce the desired temporal misalignment, enhancing sensor fusion effectiveness. The distinct hardware configurations, particularly with the magnetometer sensor being the slave to the IMU and the presence of FIFO buffers in IMU devices, contribute to this desynchronization. This scenario reveals the potential for exploiting hardware-level interactions between sensors to achieve precise temporal misalignment, thereby advancing the development of effective recognition systems. However, this also uncovers security vulnerabilities in certain smartphone models, where this misalignment could be exploited for unauthorized data extraction.
Figure 3. Normalized sensor responses to audio: Gyroscope and accelerometer readings are shown with the y-axis scaled from 0 to 1. The accelerometer exhibits a higher response compared to the gyroscope, yet both display discernible patterns.
@@ 4.3. Data collection
To enhance the performance of our STAG system in interpreting speech labels, we collected over 16,000 new sensor readings from accelerometers and gyroscopes (Author, 2024) [^1] Reference witheld due to double blind.. This was in addition to utilizing 20,000 existing readings from the StealthyIMU dataset (ste, 2024).
For textual data, we utilized Google’s Directions API to derive travel routes between a combination of over 100 cities and prominent landmarks within the United States, ensuring consistency with the StealthyIMU dataset and aiming to enhance prediction accuracy. The city list was generated using GPT and validated through the Maps API. These routes were converted into spoken instructions using Google’s Text-to-Speech service, which employs the WaveNet model—technology also utilized by Google’s voice assistant (goo, ). To extract all SLU entities from each route description, we employed SpaCy (spa, ), a Python library for natural language processing, and combined it with regular expression techniques and the GPT-4 API. This rigorous process ensures that our system accurately interprets the speech labels. All collected data, including each wave file and associated metadata, were organized and stored in CSV format.
Figure 5. Overview of the STAG Architecture and its Accuracy Assessment: Panel (a) displays the results of predictions made by the LightGBM model on normalized accelerometer data. Panel (b) illustrates the outcome of cubic spline interpolation applied to the data, enhancing resolution for improved analysis. Panel (c) demonstrates the accuracy enhancement achieved through the combination of the original and inferred accelerometer samples, with respective R-squared values indicating the goodness of fit. The overall flowchart depicts the preprocessing steps applied to accelerometer and gyroscope data, followed by feature extraction using LightGBM, interpolation, and final analysis in the SLU model.
@@ 4.5. STAG Architecture
STAG starts by misaligning accelerometer and gyroscope data sequences, optimizing them through a dual-method upsampling of LightGBM and cubic spline interpolation, with respective R-squared values shown in Figure 5(a) and (b). Data accuracy is further refined by merging the prediction and interpolation outputs, achieving an enhanced R-squared value depicted in Figure 5(c).
@@ 5. Evaluation
We present a thorough evaluation of STAG, comparing it against established methods like InertiEar (Gao et al., 2022) and StealthyIMU. We focus on assessing STAG’s performance through a structured setup involving data bifurcation, training, testing, and key metrics like WER, SER, and SEER. This detailed analysis aims to validate STAG’s effectiveness and highlight its advancements in sensor data analysis for mobile devices.
@@ 5.1.2. Training:
During this phase, the training subset undergoes preprocessing steps like noise reduction and Z-score normalization. These steps prepare the data for processing by the LightGBM algorithm. The training phase is divided into two parts. First, we train a LightGBM model with odd accelerometer values as features and even accelerometer values as feedback. This itself gives us a good baseline, as seen in Figure 5(a). After the model is trained, we carry out cubic spline interpolation on odd accelerometer values, combining the output of the interpolation with the one from the model to train a new combined model. We use mean squared error and R-squared to finetune our model.
@@ 5.3. Results
We now comparatively analyze STAG’s peformance against two established approaches in SLU applications: InertiEar (Gao et al., 2022) and StealthyIMU (Sun et al., 2023). Our discussion is anchored by the key results detailed in Table 4, showcasing STAGs’s WER of 13% and SEER of 20%. These metrics underscore the general accuracy in transcribing function words and also the STAG’s adeptness at capturing semantically significant entities such as names and locations. This performance is significantly better than that of StealthyIMU and InertiEar, suggesting an enhanced capability for dealing with complex audio data.
StealthyIMU relies solely on accelerometer data without integrating gyroscope data, limiting its ability to recognize complex speech patterns effectively. In contrast, InertiEar attempts to utilize both accelerometer and gyroscope data. However, their methodology largely involves using temporally aligned sensor data, which does not significantly enhance accuracy due to the minimal temporal disparity between the data streams. This alignment results in suboptimal performance as the interleaving of synchronized data provides no substantial improvement over using single sensor data.
STAG distinguishes itself by employing a novel method of inducing deliberate temporal misalignment between accelerometer and gyroscope data. This strategic misalignment allows STAG to exploit the unique data characteristics of each sensor type more effectively, significantly enhancing the system’s ability to decode complex audio signals. By shifting the data streams in time, STAG achieves a higher fidelity in reconstructing speech, demonstrated by superior WER and SEER results. Also, as detailed in Section 4.2, STAG leverages the high SNR of the accelerometer’s Z-axis and intelligently combines it with the X and Y axes of the gyroscope, which are temporally misaligned. This approach not only optimizes SNR but also ensures a more meaningful correlation between the data streams, enhancing both the accuracy and reliability of the sensor data utilized by STAG.
This comparative analysis conclusively underlines the superiority of STAG’s methodology. By effectively harnessing the power of temporal misalignment, STAG advances the capabilities of SLU systems, outperforming traditional methods like those employed by InertiEar and StealthyIMU. The results validate the effectiveness of STAG and highlight its potential to redefine the landscape of sensor-based SLU technologies.
@@
STAG showed remarkable improvements: a 58% reduction in sentence error rate and an 86% reduction in word error rate, compared to data sampled at 200 Hz without up-scaling.

### v3

@@ Abstract.
Privacy attack, Smartphone VUI, Eavesdropping
ccs: Security and privacy Embedded systems securityccs: Security and privacy Side-channel analysis and countermeasures
@@ 2.1.4. Spoken Language Understanding (SLU):
Paper Sensor Recognition Evaluation Metric Accuracy @ 200 Hz
@@
PitchIn (Han et al., 2017) Acc Words (constrained) Accuracy NA
Spearphone (Anand et al., 2021) Acc Digits + Words (constrained) F-measure NA
AccelEve (Ba et al., 2020) Acc Digits + Letters (constrained) Accuracy NA
@@
InertiEar (Gao et al., 2022) Acc+Gyro Digits + Letters (constrained) Accuracy 78.8%
StealthyIMU (Sun et al., 2023) Acc Spoken Language Understanding (unconstrained)
@@
STAG [Ours] Acc+Gyro Spoken Language Understanding (unconstrained)
@@ 2.2. Related Work
AccelEve (Ba et al., 2020) marks the first study exploiting loudspeaker proximity to IMUs. It employs a deep learning model (DenseNet) to reconstruct speech from accelerometer signal spectrograms, achieving a top-1 accuracy of 78% for digits and 55% for digits plus letters. Spearphone (Anand et al., 2021) utilizes altered accelerometer readings and machine learning techniques to classify gender with over 90% accuracy and identify speakers with over 80% accuracy. Additionally, Spearphone can reconstruct speech with a word detection accuracy of 67%.
@@ 3.2. Attack Scenarios
3.3. End-to-End Attack
Figure 2. An end-to-end attack scenario: a malicious app captures IMU data from a navigation app and processes it via STAG to infer sensitive user information.
An example of an end-to-end attack is depicted in Fig. 2. In this scenario, a malicious application runs covertly in the background of a navigation app, continuously capturing IMU data without the user’s knowledge. This data is subsequently processed using the STAG framework to generate a text file containing semantic frames, which are then stored on the device. These semantic frames enable the extraction of intents and entities from each VUI response. By aggregating multiple responses over time, the attacker can infer sensitive user information—such as locations, routes, and home or work addresses—posing a significant threat to user privacy.
@@ 4.1. Temporal misalignment
Although the accelerometer instances had staggered start times, both produced identical timestamps. This occurs because Android’s Sensor Service, which mediates data from the physical sensor via the Sensor HAL, distributes the same data stream, including timestamps from the system’s monotonic clock, to all requesting apps. Consequently, each accelerometer instance receives identical sensor readings. Moreover, Android does not strictly maintain a 200 Hz sampling rate, allowing the Sensor Service to throttle data to manage power and load. As a result, both instances effectively captured the same data.
Lesson Learned: Attempting to create multiple software instances of the accelerometer within an Android application to simulate a higher sampling rate is ineffective. In the Android operating system, the Sensor Service and HAL deliver sensor data in a manner that yields identical timestamps for readings from the same physical sensor, regardless of how many instances request them. This underscores inherent limitations in the OS and hardware design, illustrating that timestamps reflect system-level data handling rather than the precise moment of physical sensor capture.
@@
The cause of this synchronized timestamp phenomenon can be traced back to two main factors. Firstly, Android’s management of sensor sampling rates is inherently flexible; while a maximum rate is specifiable, the actual sampling frequency can vary and is adjusted in real-time by the operating system through the Sensor Service. This variability in the sampling intervals could result in the gyroscope and accelerometer inadvertently aligning their readings, despite our efforts to stagger their activation. Secondly, the design of IMU sensors, which encompasses both accelerometers and gyroscopes, is inherently inclined towards synchronized functioning. These sensors are typically integrated at both hardware and software levels to facilitate efficient and precise motion detection, which naturally leads to their simultaneous data sampling and reporting. The Sensor HAL, which interfaces directly with the IMU, often timestamps data from both sensors at the same time, reflecting the synchronous operation of the IMU. The Sensor Service then delivers this data to applications without altering the timestamps. A key finding was the consistent zero variance in sampling delays, indicating that the timestamps were likely being assigned by a software component (either the HAL or the Sensor Service) based on a timer, rather than reflecting the actual hardware capture time.
Lesson Learned: Introducing gyroscope and accelerometer sensors with a staggered initialization also fails to achieve desynchronized readings, as the Android OS manages sensor sampling rates flexibly through the Sensor Service and the Sensor HAL, and IMU sensors are often designed to operate synchronously. This experiment underlines the challenge of achieving temporal misalignment through software alone, due to the inherent synchronized functioning of IMU components, the HAL’s timestamping practices, and the OS’s real-time adjustments to sampling rates
4.1.3. Scenario 3: introducing magnetometer.
The distinct hardware configurations of sensors, particularly the STMicroelectronics LSM6DSM IMU (ST, 2024) found in the Samsung A31/A33 and the Bosch BMI160 IMU (BOSCH, 2024) used in the Pixel 5, involve direct connections between the IMU and the host, with the magnetometer interfacing through the IMU. Hence, the magnetometer lacks a dedicated communication channel to the host processor and instead communicates through the IMU, often via an I2C interface.. This architectural design involves a FIFO buffer within the IMU that aids in efficient high-frequency data capture and energy savings. However, it can introduce synchronization problems when integrating a Magnetometer that does not have its own FIFO buffer and whose values are stored in the FIFO of IMU—leading to data misalignments, as our research indicates. Specifically, when the magnetometer is enabled, its data is queued into the IMU’s FIFO buffer along with the accelerometer and gyroscope readings. This extra data flow disrupts the expected timing of accelerometer and gyroscope sample retrieval in the Sensor HAL. For example, the HAL might read an accelerometer sample, then unexpectedly encounter magnetometer data, followed by a gyroscope sample, altering the timing relationship between the accelerometer and gyroscope data as observed by the Sensor Service and application. The exact disruption mechanism varies by IMU and HAL implementation, often involving the HAL reading more data than anticipated or experiencing shifted interrupt timings due to the added magnetometer entries.
Lesson Learned: Integrating a magnetometer with gyroscope and accelerometer systems can induce the desired temporal misalignment, enhancing sensor fusion. The magnetometer’s subordinate role to the IMU and the presence of FIFO buffers drive this desynchronization. Adding magnetometer data to the IMU’s FIFO disrupts the expected sequence and timing of accelerometer and gyroscope data retrieval by the HAL, resulting in observable misalignment. This interaction can be leveraged to improve recognition systems, but it also exposes security vulnerabilities in some smartphone models, where the misalignment may be exploited for unauthorized data extraction.
@@
Figure 4. Normalized accelerometer and gyroscope responses to audio. The y-axis is scaled from 0 to 1. The accelerometer shows a stronger response than the gyroscope, but both sensors exhibit discernible patterns.
@@ 4.2. Establishing Correlation
4.2.1. Physical Basis for Sensor Correlation
The observed correlation between the accelerometer’s Z-axis and the gyroscope’s X and Y-axes arises from basic motion principles. When a smartphone accelerates linearly along its Z-axis (e.g., moving vertically), inertial forces induce slight rotations around the X and Y axes. Essentially, changes in linear acceleration lead to corresponding changes in angular velocity. Moreover, precise IMU design and calibration ensure that these measured correlations reflect actual motion dynamics rather than sensor crosstalk or electromagnetic noise. Consequently, gyroscope data can reliably indicate positional changes and serve as a basis for predicting future positions.
@@ 4.3. Data collection
To enhance the performance of our STAG system in interpreting speech labels, we collected over 16,000 new sensor readings from accelerometers and gyroscopes. This was in addition to utilizing 20,000 existing readings from the StealthyIMU dataset (StealthyIMU, 2024).
For textual data, we utilized Google’s Directions API to derive travel routes between a combination of over 100 cities and prominent landmarks within the United States, ensuring consistency with the StealthyIMU dataset and aiming to enhance prediction accuracy. The city list was generated using GPT and validated through the Maps API. These routes were converted into spoken instructions using Google’s Text-to-Speech service, which employs the WaveNet model—technology also utilized by Google’s voice assistant (Cloud, 2024). To extract all SLU entities from each route description, we employed SpaCy (spaCy, 2024), a Python library for natural language processing, and combined it with regular expression techniques and the GPT-4 API. This rigorous process ensures that our system accurately interprets the speech labels. All collected data, including each wave file and associated metadata, were organized and stored in CSV format.
Figure 6. Overview of the STAG architecture and its accuracy assessment. (a) LightGBM predictions on normalized accelerometer data. (b) Cubic spline interpolation results, enhancing data resolution. (c) Accuracy improvement from combining original and inferred accelerometer samples, with R-squared values indicating goodness of fit. The flowchart depicts the preprocessing, LightGBM-based feature extraction, interpolation, and final SLU model analysis.
@@ 4.5. STAG Architecture
STAG begins by intentionally misaligning the accelerometer and gyroscope data sequences. These sequences are then optimized using a dual-method upsampling approach employing both LightGBM, which predicts missing data points based on the correlation learned from the training data, and cubic spline interpolation, which smooths the data. The effectiveness of each method is assessed using their respective R-squared ($R^{2}$) values, which indicate the proportion of variance in the data explained by the model. Higher $R^{2}$ values signify a better fit to the data. These $R^{2}$ values are depicted in Figure 6(a) for LightGBM and Figure 6(b) for cubic spline interpolation. Data accuracy is further refined by merging the prediction and interpolation outputs, achieving an enhanced $R^{2}$ value depicted in Figure 6(c).
@@ 5. Evaluation
We present a thorough evaluation of STAG, comparing it against established methods like InertiEar (Gao et al., 2022) and StealthyIMU and accounting for motion artifacts. We focus on assessing STAG’s performance through a structured setup involving data bifurcation, training, testing, and key metrics like WER, SER, and SEER. This detailed analysis aims to validate STAG’s effectiveness and highlight its advancements in sensor data analysis for mobile devices.
@@ 5.1.2. Training:
During this phase, the training subset undergoes preprocessing steps like noise reduction and Z-score normalization. These steps prepare the data for processing using the LightGBM algorithm. The training phase is divided into two parts. First, we train a LightGBM model with odd accelerometer values as features and even accelerometer values as feedback. This itself gives us a good baseline, as seen in Figure 6(a). After the model is trained, we carry out cubic spline interpolation on odd accelerometer values, combining the output of the interpolation with the one from the model to train a new combined model. We use mean squared error and R-squared to finetune our model.
@@ 5.3. Results
5.3.1. Comparison with State-of-the-art
STAG’s capabilities are compared against two established SLU approaches: InertiEar (Gao et al., 2022) and StealthyIMU (Sun et al., 2023). Table 4 highlights the key performance metrics, demonstrating STAG’s superior accuracy with a WER of 13% and a SEER of 20%. These results underscore STAG’s effectiveness in transcribing functional words and capturing semantically significant entities like names and locations. Notably, STAG achieved a 58% reduction in SER and an 86% reduction in WER compared to data sampled at 200 Hz without up-scaling, as shown in Table 4.
The enhanced performance of STAG can be attributed to its novel approach of inducing temporal misalignment. StealthyIMU, which relies solely on accelerometer data, struggles to recognize complex speech patterns. InertiEar, while using both accelerometer and gyroscope data, uses temporally aligned data. The minimal temporal disparity in InertiEar’s aligned data provides no substantial improvement over single-sensor data, leading to suboptimal performance.
STAG’s strategic temporal misalignment between accelerometer and gyroscope data allows it to exploit the unique characteristics of each sensor type more effectively. As detailed in Section 4.2, STAG leverages the high SNR of the accelerometer’s Z-axis and combines it with the temporally misaligned X and Y axes of the gyroscope. This approach optimizes SNR and ensures a more meaningful correlation between data streams, enhancing the accuracy and reliability of the sensor data. Shifting the data streams in time allows STAG to achieve a higher fidelity in reconstructing speech, as evidenced by the superior WER and SEER results. This comparative analysis conclusively underlines the superiority of STAG’s methodology. By effectively harnessing the power of temporal misalignment, STAG advances the capabilities of SLU systems, outperforming traditional methods like those employed by InertiEar and StealthyIMU.
Table 5. Performance of STAG and StealthyIMU under various motion conditions.
Scenario WER (%)
StealthyIMU (Tabletop, Ideal) (Sun et al., 2023) 78.75
STAG (Tabletop, Ideal) 13.02
STAG (Handheld) 38.02
STAG (Car, Stationary) 54.77
STAG (Car, Moving) 67.58
5.3.2. Robustness to Motion Artifacts
To evaluate the robustness of STAG and its resilience to motion artifacts, we tested it under various conditions, as summarized in Table 5. The optimal scenario for STAG involves the smartphone being placed on a hard, stationary tabletop surface, where external interference vibrations are minimized. For comparison, we also evaluated the state-of-the-art (SOTA) StealthyIMU under the same conditions.
In the most common usage scenario, where the smartphone is handheld, we observed a notable increase in the WER by 25% compared to the ideal case. Additionally, we tested STAG under more challenging conditions, such as when the smartphone was mounted on a car stand, both stationary and while driving over an uneven road with speed bumps. Even under these worst-case scenarios, STAG demonstrated superior performance compared to StealthyIMU.
Table 6. End-to-end results: VUI responses, entities detected by STAG, and resulting privacy leaks.
Type Privacy Entity User Request VUI Response Entities Detected Privacy Leak
WEATHER LOCATION What’s the weather like today? Today in San Antonio you can expect a high of 87 degrees. There is currently a severe thunderstorm watch in effect. {type: ’city’, filler: ’memphis’} NO
WEATHER LOCATION Is it going to rain at Huntington Beach? Probably not today in Huntington Beach. There’s only a very slight chance of rain. {type: ’city’, filler: ’huntington beach’} YES
WEATHER LOCATION Is there a chance of snow in Chandler? Tonight in Chandler there’s no chance of snow. {type: ’city’, filler: ’chandler’} YES
REMINDER TODO/TIME Add a reminder to pick up prescription OK, pick up prescription. When do you want to be reminded? {type: ’todo’, filler: ’pick up; party; supplies’} NO
REMINDER TODO/TIME Add a reminder to buy toothpaste OK, buy toothpaste. When do you want to be reminded? {type: ’todo’, filler: ’buy; toothpaste’} YES
REMINDER TODO/TIME Remind me to do dishes tomorrow at 5 PM All right, I’ll remind you tomorrow at 5 PM. {type: ’time’, filler: ’tomorrow; at; five; pm’} YES
STOCK UPDATES SEARCH What is the stock price of General Electric? General Electric Company closed down 3.4% on Friday at \$81.0. {type: ’company’, filler: ’general motors; company’} NO
STOCK UPDATES SEARCH What is the stock price of Berkshire Hathaway? Berkshire Hathaway Class A closed up 0.75% on Friday at \$432.00. {type: ’company’, filler: ’berkshire hathaway’} YES
SUNSET/SUNRISE LOCATION When did sunset in Baltimore? In Baltimore today, the sunset at 4:55 PM. {type: ’city’, filler: ’baltimore’} YES
SUNSET/SUNRISE LOCATION What is the sunrise in Buffalo? In Buffalo today, sunrise is at 7:01 AM. {type: ’city’, filler: ’tampa’} NO
AQI INDEX LOCATION AQI in Glendale According to Purple Air, the air quality near the center of Glendale is good with an index ranging from 18 to 26. {type: ’city’, filler: ’chicago’} NO
AQI INDEX LOCATION AQI in Amarillo According to Air Now, the air quality near the center of Amarillo is good with an index of 21. {type: ’city’, filler: ’amarillo’} YES
5.3.3. End-to-End Attack Scenario
We conducted a series of experiments simulating real-world usage to demonstrate STAG’s capability in a realistic end-to-end attack scenario. In our setup, a user interacts with a voice assistant via common VUI requests while STAG, running in the background, captures and processes IMU sensor data. This mirrors a real-world attack where a malicious app could exploit zero-permission sensor access to eavesdrop on user interactions. The experiments focused on common VUI interactions related to weather, reminders, stock updates, sunset/sunrise times, and air quality index (AQI).
For each VUI request type, we present the specific user request, the corresponding VUI response, the entities detected by STAG, and whether a privacy leak occurred. As shown in Table 6, STAG successfully identified most of the target privacy entities. These results highlight the potential for STAG to compromise user privacy by exploiting seemingly innocuous VUI interactions.
Table 7. Prevention strategies for STAG-based attacks.
Prevention Strategy Responsible Party Practicality
Stricter Access Controls Operating System Developers High – Integrates with existing permission systems with minimal additional overhead.
Noise Injection Techniques Application and OS Developers Medium – Requires careful implementation to maintain application performance while obfuscating data.
Direct Host Connection for Sensors Hardware Manufacturers Low to Medium – May necessitate hardware redesign, increasing development time and costs.